LINE INTENSITIES FOR THE 10 μm BANDS OF SO_{2}

S. J. WETZEL and \underline{P. M. CHU}, Analytical Chemistry Division, CSTL, National Institute of Standards and Technology, Gaithersburg, MD 20899, USA; W. J. LAFFERTY, Optical Technology Division, Physics Laboratory, National Institute of Standards and Technology, Gaithersburg, MD 20899, USA; A. PERRIN and J. -M. FLAUD, Laboratoire de Photophysique Moleculaire, CNRS, Universite Paris Sud, Bat 210, Campus D’Orsay, F–91405 ORSAY, Cedex, France; PH. ARCAS and G. GUELACHVILI, Laboratoire de Physique Moleculaire et Applications, CNRS, Universite Paris Sud, Bat 350, Campus D’Orsay, F–91405 ORSAY, Cedex, France; L. BROWN, Jet Propulsion Laboratory, Pasadena, CA 91109, USA.

Using both high resolution (R = 0.003 cm^{−1}) and medium resolution (R = 0.12 cm^{−1}) Fourier transform spectra recorded at LPMA Orsay and Kitt Peak, and at NIST Gaithersburg respectively, it has been possible to measure a large set of individual line intensities for the ν_{1} and ν_{3} bands of SO_{2} in the 950—1350 cm^{−1} spectral region. These intensities were introduced in a least squares fit calculation allowing one to get the expansion of the transition moment operator of the ν_{1} and ν_{3} bands of SO_{2}. For these intensities calculations, the theoretical model takes into account the vibration-rotation interactions linking the upper levels involved in the ν_{1}, 2ν_{2} and ν_{3} interacting bands of SO_{2}. Finally a synthetic spectrum of the 10 μm bands of SO_{2} has been generated, using for the line intensities the dipole moment expansion determined in this work and for the line positions the parameters and the Hamiltonian matrix given in a previous analysis [J.-M. Flaud, A. Perrin, L. M. Salah, W. J. Lafferty and G. Guelachvili, J. Mol. Spectrosc. 160, 272—278 (1993)].